Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ (a) mean,​ (b) median,​ (c) mode, and​ (d) midrange for the given sample data in millions of dollars. Given that these are the top 10​ salaries, do we know anything about the salaries of TV personalities in​ general? Are such top 10 lists valuable for gaining insight into the larger​ population? 37.6   35.4   34   26.2   15.9   14   11.3   10.2   9.6   9 a. The mean is nothing. ​(Type an integer or a​ decimal.) b. The median is nothing. ​(Type an integer or a​ decimal.) c. Select the correct choice below and fill in any answer boxes in your choice. A. The mode is nothing. ​(Use a comma to separate answers as​ needed.) B. There is no mode. d. The midrange is nothing. ​(Type an integer or a​ decimal.) Given that these are the top 10​ salaries, do we know anything about the salaries of TV personalities in​ general? A. Since the​ mean, median, and midrange are based on a small​ sample, no information is given on the salaries of TV personalities in general. B. Since the sample values are the 10​ highest, they give almost no information about the salaries of TV personalities in general. C. Since the sample values give information about one segment of the salaries of TV​ personalities, they give a lot of information about the salaries of TV personalities in general. D. Since the​ mean, median, and midrange are relatively reliable even with small​ samples, a lot of information is given on the salaries of TV personalities in general. Are such top 10 lists valuable for gaining insight into the larger​ population? A. ​No, because such top 10 lists represent an extreme subset of the population rather than the larger population B. ​No, because the​ mean, median, and midrange are based on a small sample C. ​Yes, because the​ mean, median, and midrange are relatively reliable even with small samples D. ​Yes, because such top 10 lists give partial information about the population
Answer #1

Answer:
a) The mean is the usual average, so I'll add up and then divide:

= (37.6,35.4,34,26.2,15.9,14,11.3,10.2,9.6,9)/10

=20.32

The mean is 20.32

b) The median is the middle value. In a list of ten values, that will be the (10 + 1) ÷ 2 = 5.5-th value; the formula is reminding me, with that "point-five", that I'll need to average the fifth and sixth numbers to find the median. The fifth and sixth numbers are the last 15.9 and the first 14, so:

The median is 14.95

c) The mode is the number that is repeated most often, but all the numbers in this list appear only once, so there is no mode

d)The formula to find the midrange = (maximum + minimum) / 2.

Midrange = (37.6+9)/2=23.3

The mid range is 23.3
Since the mean, median, and midrange are based on a small sample, no information is given on the salaries of TV personalities in general.

A.No, because such top 10 lists represent an extreme subset of the population rather than the larger population.